### v1

# namerefThe definition of Are magnetic fields in cosmic voids primordial?

Abstract
The nature of magnetic fields in the voids of the large-scale structure of the Universe has been a multifaceted open puzzle for decades. On one hand, their origin is not clear with most of the magnetogenesis models using physics beyond the standard model in the early Universe, and on the other hand, their existence and potential role in explaining the spectra of TeV blazars have been intensely debated in the past decade. Here, we show that void fields must exist and can be generated using classical electrodynamics in the late Universe. Specifically, we use the dipole component of the galactic fields, decaying as , and show that this generates space-filling magnetic fields in voids with sufficient amplitude to explain the blazar spectra and the lack of GeV halos observed by Fermi-LAT. Our results also show that the void fields might be too strong even for Cherenkov Telescope Array to observe the GeV halos. While primordial magnetic fields could still have been generated through some non-standard physics in the early Universe, our results show that the need for primordial magnetic fields is not strengthened by the presence of void magnetic fields.
I Introduction
The standard model of cosmology (CDM) has been very successful in explaining a wide range of observations of the Universe and its history, including Hubble’s law of the expansion of space [1], numerous features of the cosmic microwave background [2], and the large-scale structure (LSS) of the Universe. However, there are still many open questions related to, for example, the inflationary stage at the beginning [3, 4], cosmological phase transitions that may have occurred later [5, 6], the asymmetry between matter and antimatter [7], nature of dark matter and dark energy [8], relative abundance of light elements [9], Hubble tension [10], tension [11], galaxy formation [12], and galactic magnetism [13, 14].
It has been shown that magnetic fields can have an impact on several of the aforementioned phenomena and can, thus, help shape the Universe significantly (for reviews, see Refs. [15, 16, 17]). Some of the leading theories of inflation [e.g. 18] predict the production of magnetic fields that can survive until today in the voids of LSS [19, 20]. These primordial fields would also impact the spectrum of the relic gravitational waves [21] from inflation [22] as well as from (first-order) cosmological phase transitions [23]. If primordial magnetic fields are helical as predicted by some inflationary magnetogenesis models, they can even help explain the asymmetry between matter (baryons) and antimatter (antibaryons) through the conservation of the Chern-Simons number [24, 25]. If present, primordial magnetic fields have also been shown to play a role in many important tensions in modern cosmology, like the Hubble tension and the tension [26, 27, 28], and the lithium abundance problem [29]. Beyond their direct effect on the physics of the early Universe, primordial magnetic fields can influence how cosmic structures from in the post-recombination Universe. Through additional density fluctuations in the primordial plasma, they modify the seeds of cosmic structure formation, which affects galaxy formation [30] and the reionization epoch in the late Universe, and therefore the global 21cm signal [31, 32, 33].
Primordial magnetic fields are important not only for astrophysics, but they are also relevant to fundamental physics at the microscopic levels. One of our leading ways to probe fundamental high-energy physics is via particle accelerators, which can reach energies up to TeV scales. However, the standard model of the particle physics (SM) may well remain valid for many orders of magnitude above the current particle accelerators, which will not be accessible by accelerators in the foreseeable future [34]. Consequently, exploring the physics beyond the SM (BSM) remains an uncertain task with accelerator physics. On the other hand, the early Universe can help us get insight into some of these questions, since the energy was much higher than accelerator energies (maybe even up to the Planck mass TeV). Since all the magnetogenesis models proposed in the literature employ BSM physics, a detection of these primordial fields would open a direct window into the extension of the SM.
With such high significance of primordial magnetic fields for addressing critical questions studied across different communities, there are several ongoing efforts to detect them (for a review, see [35]). However, the only place in the present-day Universe where these primordial fields can be expected to be in their original form are in the voids of the LSS. In the other regions of the Universe, the signature of the primordial magnetic field is most likely contaminated by astrophysical processes in the late Universe. One of the main such processes is the dynamo mechanism, which amplifies magnetic fields exponentially by randomly stretching, twisting and folding magnetic field lines through turbulent plasma motion [36, 37, 38]. This has also been demonstrated in cosmological simulations [39, 40, 41]. These highly non-linear plasma processes make it practically impossible to deduce the initial conditions of the magnetic fields before the emergence of cosmic structures, leaving the void fields as the only potential relic of the primordial field.
While theorized for decades, the first direct evidence for the existence of void fields was reported in 2010 [42, 43], providing a proof of concept of the viability of primordial fields as a probe of fundamental physics. There, the authors used the spectra of TeV blazars to deduce the lower limit of the field strength of void magnetic fields of G for coherence length scales above 1 Mpc. The high volume-filling factor required of void fields for these observations makes any astrophysical process disfavored over a primordial origin, which produces large scale fields throughout the Universe and, hence, also in voids [15, 16, 17]. However, other plasma processes have also been proposed to explain the observed TeV blazar spectra [44, 45]. Given the extreme range of density and energy scales involved, both analytical and numerical studies have not yet provided a conclusive resolution, and the existence of void fields remains an open problem. A clear proof of the existence of void fields would be the observation of GeV halos around TeV blazars, which would allow for a direct measurement of the void field strength.
In this paper, we propose that void magnetic fields need to exist, resulting solely from classical electrodynamics in the late Universe. Specifically, we show that the dipole component of the galactic magnetic fields can produce space-filling void fields of the magnitude within the range suggested by the blazar spectra.
II Simple analytical estimate
Let us consider spiral galaxies, like our Milky Way, surrounding a void of the LSS. These galaxies will have a magnetic field with a dipole component that decays as . The sum of all of the dipole fields of the galaxies will result in an extragalactic magnetic field (EGMF) with 100% volume-filling factor. We first estimate the root mean square (RMS) of this EGMF at the center of a large void, where it will be the weakest. In the next section, we use numerical tools to calculate the characteristic fields in a void and the associated power spectrum.
Let us consider a large void. For our order of magnitude analysis in this section, we assume a spherical void of the radius Mpc. A void is generally surrounded by filaments and nodes, but for this crude estimate, let us consider galaxies to be distributed uniformly over the shell of the 20 Mpc sphere. A more realistic distribution of galaxies should provide approx. corrections and should only strengthen the fields at the center since the fields will cancel out most effectively for a uniform distribution. For a typical intergalactic distance of Mpc, we can easily obtain the number of galaxies surrounding the void
where is the fraction of all galaxies that have appreciable magnetic dipoles. For now, we can crudely take all the dipoles to have a similar strength of G with . However, this could later be changed to incorporate some distribution of the dipole strength across galaxies. The dipole of each galaxy decays as to contribute from one galaxy at the center of the void, given by
Here, is the distance from the center of the galaxy at which G. According to Refs. [46, 47, 48], the out-of-plane component of the galactic magnetic field remains of the order of G up to the distance of kpc, so we set kpc. We then have galaxy dipoles each contributing a magnetic field of the strength at the center of the void. In this simple estimate, we neglect possible correlations of the galactic dipole fields and assume a random orientation that is uniformly distributed across the solid angle. We can use the Box-Muller transformation [49] to find that each component of will follow a Gaussian distribution normalized to a fixed radius. Upon adding of , each component of the resulting field will be a normally distributed variable. The magnitude of this total magnetic field at the center will have the Maxwell-Boltzmann distribution with the RMS given by . We have also confirmed this numerically with Mathematica [50]. Combining all of this, we obtain a rough estimate for the resulting field
This is an estimate for the RMS value of the magnetic field at the center of the void. Note that the RMS field (3c) will be minimal at the center of the void. In the next section, we calculate the RMS value of the magnetic field numerically in the entire void and estimate its power spectrum.
III Numerical analysis
In order to calculate the magnetic field in the whole void numerically, we start with the relatively simple configuration assumed in the previous section, i.e., of a totally empty void with galaxies uniformly distributed over its surface. For simplicity, we consider a cubic void rather than a spherical one in the numerical analysis, with the assumption that the void is sufficiently large that its shape does not impact the magnetic fields appreciably in most of the void volume.
The void is modeled as a cube with galaxies spread on the surface of the cube as point particles. The number of galaxies is taken to be 504, which is approximately the number of galaxies expected around a spherical void with 20 Mpc, Mpc and . Each galaxy is modeled as a randomly oriented dipole with a constant strength G at kpc.
The power spectrum of the resulting fields, defined as
where | |||
turns out to be that of a white noise, , which is expected given the lack of any correlations imposed on the system. The characteristic magnetic field at the scale of 1 Mpc, defined as
for , is around G, about two and a half orders of magnitude stronger than the rough estimate of the field at the center of the void (3c). The characteristic field across length scales is shown in Fig. 1. Also, this enhancement is not surprising as Eq. 3c gives the minimal field in the void, while Eq. 6 gives the mean field on the scale of 1 Mpc.
IV Conclusion
With our brief estimate, we have shown that it is indeed possible to generate void magnetic fields strong enough to explain the observed blazar spectra without resorting to physics beyond the SM. We use the dipole component of galactic fields to show that they should pervade throughout the voids, generating fields well above the lower bounds set by the blazar spectra.
While magnetic fields could still have been generated in the early Universe (possibly through non-standard physics), our results show that the need for primordial magnetic fields is not strengthened by the presence of void magnetic fields.